OpenStudy (anonymous):

Explain why this problem is in a number theory book? I feel like it should be in a Cal 3 book. Prove that if (xₒ,yₒ) is a solution of the linear Diopahntine equation ax - by = 1, then the area of the triangle whose vertices are (0,0), (b,a) and (xₒ,yₒ) is 1/2? This could easily be proven using cross product

OpenStudy (usukidoll):

The Diophantine equation was in my number theory book as well. An equation which only integer solutions are allowed.

OpenStudy (anonymous):

eh, probably because it mentions linear Diophantine equations?

OpenStudy (anonymous):

And there is this problem too. A man pays $1.43 for some apples and pears. If pears cost 17 cents each, and apples, 15 cents each, how many of each did he buy? O.O what the heck? this is 5th grade problem!

OpenStudy (usukidoll):

oh I had that problem too... hey are you reading Number Theory by George Andrews?

OpenStudy (anonymous):

@UsukiDoll yep :)

OpenStudy (anonymous):

anyways, this just follows from the determinant: $$\left|\det\begin{bmatrix}b&x_0\\a&y_0\end{bmatrix}\right|=\left|by_0-ax_0\right|=1$$ which is the area of the whose edges are given by \((b,a),(x_0,y_0)\), so the triangle is simply half that: \(1/2\)

OpenStudy (usukidoll):

I've used that book last year in my Elementary Number Theory class XDDDDDDDDDDD!!!!!!!! Luckily I didn't have to do that problem, but I was stunned when I saw that.

OpenStudy (usukidoll):

I've heard that it's the most popular book.

ganeshie8 (ganeshie8):

Mostly the authors choose trivial examples in the start because the focus is on succcessfully applying a particular method, not on getting a quick answer

ganeshie8 (ganeshie8):

the second question translates to solving linear diophantine eqn \[17a+15p = 143\]

OpenStudy (anonymous):

that problem is fine, it's a Diophantine equation, too: $$15a + 17p = 143\\ 15a=143\pmod{17}\implies -2a=-10\pmod{17}\implies a=5\pmod{17}\\17p=143\pmod{15}\implies 2p=10\pmod{17}\implies p=5\pmod{15}$$

OpenStudy (anonymous):

oops, the second line is broken

OpenStudy (anonymous):

$$2p=8\pmod{15}\implies p=8\pmod{15}$$

OpenStudy (anonymous):

O: I did the second problem by guess and check. Never mind about it being 5th a grade problem ^^

OpenStudy (anonymous):

I need someone to confirm this problem. Hold on...